Almost all medical schools in the United States require students to take the Medical College Admission Test (MCAT). The exam is composed of three multiple-choice sections (Physical Sciences, Verbal Reasoning, and Biological Sciences). The score on each section is converted to a 15-point scale so that the total score has a maximum value of 45. The total scores follow a Normal distribution, and in a certain year the mean was 24.7 with a standard deviation of 7.3. There is little change in the distribution of scores from year to year. (Round your answers to four decimal places.)

(a) What proportion of students taking the MCAT had a score over 30?

(b) What proportion had a score between 20 and 24?
Answer #1
mean μ= 24.7
standard deviation σ= 7.3000
for normal distribution z score =(X-μ)/σ

a)

probability =P(X>30)=P(Z>(30-24.7)/7.3)=P(Z>0.73)=1-P(Z<0.73)=1-0.7673=0.2327

b)

probability =P(20<X<24)=P((20-24.7)/7.3)<Z<(24-24.7)/7.3)=P(-0.64<Z<-0.1)=0.4602-0.2611=0.1991